# Consensus over Clustered Networks Using Intermittent and Asynchronous Output Feedback

### v2

@@ VI Robustness
where $\delta_{p},\delta_{pr}\in\mathbb{R}^{m}$ model non-vanishing measurement noise. Typically, measurement noise is additively injected, e.g., $y_{p}+\delta_{p}$, which would imply $\eta_{p}^{+}=\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(y_{q}-y_{p}+\delta_{q}-\delta_{p})$. Nevertheless, we consider the lumped perturbation $\delta_{p}$ instead of $\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(\delta_{q}-\delta_{p})$ for convenience. A similar statement follows for $\delta_{pr}$. Nearly all other components of $\mathcal{H}$, that is, the flow equations of $x^{\circ}$, $\tilde{\eta}$, $\tilde{\zeta}$, $\tau$, and $\rho$, the definitions in (11) and (12), and the jump equations/inclusions of $x^{\circ}$, $\tau$, and $\rho$, will remain unchanged. Yet, under the jump equations of (52) and (53), the jump equations of $\tilde{\eta}_{p}$ and $\tilde{\zeta}_{pr}$ do change.
@@
Let $\mathcal{D}\subseteq\mathbb{R}^{mN}\times\mathbb{R}^{mNM^{\ast}}$, and consider the hybrid system $\mathcal{K}$ with flow set $\widetilde{C}\triangleq C\times\mathcal{D}$, jump set $\widetilde{D}\triangleq D\times\mathcal{D}$, single-valued flow map $\tilde{f}\colon\mathcal{X}\times\mathcal{D}\to\mathcal{X}$ by $\tilde{f}(\xi,\delta)=f(\xi)$, and set-valued jump map $\widetilde{G}\colon\mathcal{X}\times\mathcal{D}\rightrightarrows\mathcal{X}$. Furthermore, let $\delta^{r}\triangleq(\delta_{pr})_{p\in\mathcal{V}}$, $\delta\triangleq((\delta_{p})_{p\in\mathcal{V}},(\delta^{r})_{r\in[M^{\ast}]})\in\mathcal{D}$,
$\widetilde{G}(\xi,\delta)$$\triangleq\{\widetilde{G}_{p}(\xi,\delta)\colon(\xi,\delta)\in D_{p}\times\mathcal{D}\text{ for }p\in\mathcal{V}\}$
$\phantom{\triangleq}\ \cup\{\widetilde{G}^{r}(\xi,\delta)\colon(\xi,\delta)\in D^{r}\times\mathcal{D}\text{ for }r\in[M^{\ast}]\},$
$\widetilde{G}_{p}(\xi,\delta)$ $\triangleq\left[\begin{array}[]{c}x^{\circ}\\ \left[\tilde{\eta}_{1}^{\top},...,\tilde{\eta}_{p-1}^{\top},\delta_{p}^{\top},\tilde{\eta}_{p+1}^{\top},...,\tilde{\eta}_{N}^{\top}\right]^{\top}\\ \tilde{\zeta}\\ \left[\tau_{1},...,\tau_{p-1},[T_{1}^{p},T_{2}^{p}],\tau_{p+1},...,\tau_{N}\right]^{\top}\\ \rho\end{array}\right],$ (54)
$\widetilde{G}^{r}(\xi,\delta)$ $\triangleq\left[\begin{array}[]{c}x^{\circ}\\ \tilde{\eta}\\ \left[\tilde{\zeta}_{1}^{\top},...,\tilde{\zeta}_{r-1}^{\top},\big{(}\tilde{\zeta}_{pr}^{+}\big{)}_{p\in\mathcal{V}}^{\top},\tilde{\zeta}_{r+1}^{\top},...,\tilde{\zeta}_{M^{\ast}}^{\top}\right]^{\top}\\ \tau\\ \left[\rho_{1},...,\rho_{r-1},[T_{3}^{r},T_{4}^{r}],\rho_{r+1},...,\rho_{M^{\ast}}\right]^{\top}\end{array}\right],$
@@
Theorem 2.
If there exist timer parameters $0<T_{1}^{p}\leq T_{2}^{p}$ for every $p\in\mathcal{V}$ and $0<T_{3}^{r}\leq T_{4}^{r}$ for every $r\in[M^{\ast}]$, a constant $\sigma$, controller matrix $K_{u}$, estimator matrices $K_{\eta,p}$ for each $p\in\mathcal{V}$, estimator matrices $K_{\zeta,pr}$ for every $(p,r)\in\mathcal{V}\times[M^{\ast}]$, and symmetric, positive definite matrices $P_{1}$, $P_{2}(\tau)$, and $P_{3}(\rho)$ that satisfy the inequality $\mathsf{M}(\tau,\rho)<\mathbf{0}$ for all $(\tau,\rho)\in\mathcal{T}\times\mathcal{R}$, then the set $\mathcal{A}$ in (27) is ISS for the hybrid system $\mathcal{K}$ with data $(\widetilde{C},\tilde{f},\widetilde{D},\widetilde{G})$. In particular, for every solution pair $(\phi,\delta)$ of $\mathcal{K}$,
@@
where $N^{\ast}\triangleq N+M^{\ast}\in\mathbb{Z}_{>0}$, [^8] The matrices $P_{2}(\tau)$ and $P_{3}(\rho)$ are defined in (30), while the definitions of the constant vectors $T_{2}$ and $T_{4}$ are located above (30).
$p_{\max}$ $\triangleq\lambda_{\max}(\text{diag}(P_{2}(T_{2}),P_{3}(T_{4})))\in\mathbb{R}_{>0},\text{ and }$
$\kappa_{2}$ $\triangleq\sqrt{\frac{p_{\max}N^{\ast}(2-\exp(-\mu T_{\min}/\alpha_{2}))}{\alpha_{1}(1-\exp(-\mu T_{\min}/\alpha_{2}))}}\in\mathbb{R}_{>0}.$
@@
$=\delta_{q}^{\top}P_{2,q}\exp(\sigma\tau_{q}^{+})\delta_{q}+(\delta^{s})^{\top}P_{3,s}\exp(\sigma\rho_{s}^{+})\delta^{s}$
$\phantom{=}\ -\tilde{\eta}_{q}^{\top}P_{2,q}\tilde{\eta}_{q}-\tilde{\zeta}_{s}^{\top}P_{3,s}\tilde{\zeta}_{s}$
@@
$\Delta V(\xi)$ $\leq((\delta_{p})_{p\in\mathcal{V}})^{\top}P_{2}(T_{2})(\delta_{p})_{p\in\mathcal{V}}$ (57)
$\phantom{\leq}+((\delta^{r})_{r\in[M^{\ast}]})^{\top}P_{3}(T_{4})(\delta^{r})_{r\in[M^{\ast}]}$
$\leq p_{\max}|\delta|_{\infty}^{2}.$
@@
$\text{dom}(\phi,\delta)\bigcap([0,t_{j}]\times\{0,1,...,j-1\})$
$=$ $\bigcup_{k=1}^{j}([t_{k-1},t_{k}]\times\{k-1\}).$
@@
$V(\phi(t,j))$ $\leq\exp\left(-\frac{\mu}{\alpha_{2}}t\right)V(\phi(0,0))$ (61)
$\phantom{\leq}+p_{\max}|\delta|_{\infty}^{2}\sum_{s=1}^{j}\exp\left(-\frac{\mu}{\alpha_{2}}(t_{j}-t_{s})\right),$
@@
Let $j^{\ast}\triangleq\lfloor j/N^{\ast}\rfloor\in\mathbb{Z}_{\geq 0}$. Note $j^{\ast}N^{\ast}\leq j$ and $j-j^{\ast}N^{\ast}\leq N^{\ast}$. The summation in (61) can be rewritten as
$\sum_{s=1}^{j}\exp\left(\!-\frac{\mu}{\alpha_{2}}(t_{j}-t_{s})\!\right)$ $=\!\sum_{s=0}^{j^{\ast}-1}\sum_{k=1}^{N^{\ast}}\exp\left(\!-\frac{\mu}{\alpha_{2}}(t_{j}-t_{sN^{\ast}+k})\!\right)$ (62)
$\phantom{=}+\sum_{s=j^{\ast}N^{\ast}+1}^{j}\exp\left(\!-\frac{\mu}{\alpha_{2}}(t_{j}-t_{s})\!\right).$
@@
where $t_{j}^{s+1}\triangleq t_{j}-t_{(s+1)N^{\ast}}$. Furthermore, since $\exp(-\tfrac{\mu}{\alpha_{2}}(t_{j}-t_{s}))\leq 1$ provided $t_{j}-t_{s}\geq 0$ and $t_{j^{\ast}N^{\ast}+1}\leq t_{j^{\ast}N^{\ast}+2}\leq...\leq t_{j}$, one has that
@@
$\sum_{s=0}^{\infty}N^{\ast}\exp\left(\!-s\frac{\mu}{\alpha_{2}}T_{\min}\!\right)$ $=N^{\ast}\sum_{s=0}^{\infty}\exp\left(\!-\frac{\mu}{\alpha_{2}}T_{\min}\!\right)^{s}$ (67)
$=\frac{N^{\ast}}{1-\exp(-\mu T_{\min}/\alpha_{2})}.$
@@
$V(\phi(t,j))$ $\leq\exp\left(-\frac{\mu}{\alpha_{2}}t\right)V(\phi(0,0))$ (69)
$\phantom{\leq}+p_{\max}|\delta|_{\infty}^{2}N^{\ast}\frac{2-\exp(-\mu T_{\min}/\alpha_{2})}{1-\exp(-\mu T_{\min}/\alpha_{2})}.$
@@
VII Simulation Example & Discussion
@@
For each $p\in\mathcal{V}$, suppose the center of mass (CM) of marine craft $p$ is fixed. The position of the CM of agent $p$ is denoted by $c_{p}\triangleq(c_{1,p},c_{2,p})\in\mathbb{R}^{2}$ and expressed in a common inertial reference frame. The heading and angular velocity are denoted by $(\theta_{p},\omega_{p})\in(-\pi,\pi]\times\mathbb{R}$, and the linear speed is denoted by $v_{p}\in\mathbb{R}$. The control input of agent $p$ is $\mu_{p}\triangleq(v_{p},\omega_{p})\in\mathbb{R}^{2}$.
For each marine craft $p\in\mathcal{V}$, let $\ell\in\mathbb{R}_{>0}$ denote the length between the CM and bow. Using the CM position $c_{p}$, heading $\theta_{p}$, and length $\ell$, the position of the bow of agent $p$ is given by $b_{p}\triangleq c_{p}+\ell(\cos(\theta_{p}),\sin(\theta_{p}))$. Figure 2 provides a clarifying illustration. The goal of this simulation example is to achieve rendezvous in the bow positions, that is, $b_{p}=b_{q}$ for all $p,q\in\mathcal{V}$. To enable the use of our consensus result, we introduce a transformation. Utilizing (70), the time derivative of $b_{p}$ is
$\dot{b}_{p}=\underbrace{\begin{bmatrix}\cos(\theta_{p})&-\sin(\theta_{p})\\ \sin(\theta_{p})&\phantom{-}\cos(\theta_{p})\end{bmatrix}}_{\triangleq R(\theta_{p})}\underbrace{\begin{bmatrix}1&0\\ 0&\ell\end{bmatrix}}_{\triangleq D}\mu_{p}.$ (71)
Since $R(\theta_{p})$ is a rotation matrix and $\ell>0$, $R(\theta_{p})^{-1}$ and $D^{-1}$ are always well-defined. Furthermore, for any velocity $\dot{b}_{p}$, the corresponding control input is given by $\mu_{p}=D^{-1}R(\theta_{p})^{-1}\dot{b}_{p}$. Given a continuously differentiable reference trajectory for the bow position $b_{p}$, one can compute the corresponding control input for (70), pointwise in time, by utilizing the inverse map $\mathsf{T}^{-1}(\theta_{p},s)\triangleq D^{-1}R(\theta_{p})^{-1}s$. Next, we provide a method for producing a reference trajectory for each $b_{p}$.
@@
where $z_{p}\in\mathbb{R}^{2}$ denotes a virtual position, $u_{p}\in\mathbb{R}^{2}$ is a virtual control input, $\mathtt{M}\in\mathbb{R}^{2\times 2}$ is a positive definite mass matrix, and $\mathtt{C}\in\mathbb{R}^{2\times 2}$ is a non-zero viscous damping matrix. The second-order ODE in (72) can be expressed as an LTI system by using $x_{1,p}\triangleq z_{p}$, $x_{2,p}\triangleq\dot{z}_{p}$, $x_{p}\triangleq(x_{1,p},x_{2,p})$, and
$A\triangleq\begin{bmatrix}0_{2\times 2}&I_{2}\\ 0_{2\times 2}&-\mathtt{M}^{-1}\mathtt{C}\end{bmatrix},\ B\triangleq\begin{bmatrix}0_{2\times 2}\\ \mathtt{M}^{-1}\end{bmatrix},\ H\triangleq\begin{bmatrix}I_{2}&0_{2\times 2}\end{bmatrix}.$ (73)
@@
By driving the auxiliary variables in $\{x_{p}\}_{p\in\mathcal{V}}$ into consensus, which will be achieved using the result from Section IV, and applying the transformation $\mathsf{T}^{-1}$, pointwise in continuous-time for each agent $p\in\mathcal{V}$, the marine crafts can accomplish rendezvous with respect to the bow positions. Two simulation cases are considered, namely, a nominal simulation that is free of perturbations and a more realistic simulation that contains measurement noise. The simulation parameters are $N=14$, $n=4$, $d=m=2$, $M=5$, $M^{\ast}=8$, $\ell=0.53$, $\sigma=500$, $\mathtt{M}=I_{2}$, $\mathtt{C}=2.66\cdot I_{2}$, $K_{u}=0.5\cdot I_{2}$, and, for all $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, $T_{1}^{p}=T_{3}^{r}=0.001$, $T_{2}^{p}=T_{4}^{r}=0.01$,
@@
Moreover, $x_{1,p}(0,0)=b_{p}(0,0)$ for each $p\in\mathcal{V}$. Under these parameters, MATLAB and the CVX Toolbox [39, 40] were employed to solve the matrix inequality $\mathsf{M}(\tau,\rho)\leq\mathbf{0}$ for points $(\tau,\rho)=(0_{N},0_{M^{\ast}})$ and $(\tau,\rho)=(T_{2},T_{4})$. Consequently, Corollary 1 implies the sufficient conditions of Theorem 1 are satisfied. Unfortunately, we are unable to provide the solution to the matrix inequality since $P(0_{N},0_{M^{\ast}})\in\mathbb{R}^{304\times 304}$ in our application.
The parameters and communication topology, i.e., the clustered network depicted in Figure 3, are identical for the nominal and perturbed simulation cases. Figures 46 correspond to the simulation results for the nominal case, while Figures 78 correspond to the simulation results for the perturbed case. To examine the robustness of our consensus controller to various perturbation magnitudes and communication graphs, a modest Monte Carlo simulation was performed with results provided in Figure 9. Note only the graph $\mathcal{G}$, perturbation magnitudes, and controller matrix $K_{u}$ were changed in the Monte Carlo simulation. The matrix $K_{u}=1.5\cdot I_{2}$ was fixed for all $30$ runs of the Monte Carlo simulation. However, the graph $\mathcal{G}$ and perturbation magnitudes were randomly selected from the set of connected graphs on $N$ nodes and the interval $[0,0.075]$ using a uniform distribution, respectively. The partition of each graph’s node set $\mathcal{V}$ was also randomly selected.
@@
VII-A Consensus under Ideal Conditions
As validated by Figures 46, the proposed control strategy in Section IV renders the set $\mathcal{A}$ globally exponentially stable under ideal conditions (i.e., under the absence of measurement noise). Consequently, the reference trajectories for the marine crafts can be synchronized exponentially fast while using intermittent and asynchronous output feedback (that is, position measurements). Also, the map $(\theta_{p},x_{2,p})\mapsto D^{-1}R(\theta_{p})^{-1}x_{2,p}$ enabled each agent $p\in\mathcal{V}$ to track their reference trajectory. Recall $x_{2,p}$ denotes the velocity of the reference trajectory for agent $p$. Therefore, we were able to achieve exponentially fast rendezvous for marine crafts modeled by (70).
Of course, driving $|\xi|_{\mathcal{A}}$ to 00 in practice would probably be undesirable as the user may not wish for potentially expensive marine crafts to crash into each other. Thus, achieving $|\xi|_{\mathcal{A}}\leq\omega$ for some appropriate $\omega>0$ may suffice instead. Given such a scenario, $|\xi|_{\mathcal{A}}\leq\omega$, (34), and (41) facilitate the calculation of a conservative, albeit finite, continuous time for achieving $\omega$-approximate rendezvous (see [41, Defintion 2]). In particular,
[…]

### v3

@@ 6 Robustness
where $\delta_{p},\delta_{pr}\in\mathbb{R}^{m}$ model non-vanishing measurement noise. Typically, measurement noise is additively injected, e.g., $y_{p}+\delta_{p}$, which would imply $\eta_{p}^{+}=\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(y_{q}-y_{p}+\delta_{q}-\delta_{p})$. Nevertheless, with a slight abuse of notation, we consider the lumped perturbation $\delta_{p}$ instead of $\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(\delta_{q}-\delta_{p})$ for convenience. A similar statement follows for $\delta_{pr}$. Nearly all other components of $\mathcal{H}$, that is, the flow equations of $x^{\circ}$, $\tilde{\eta}$, $\tilde{\zeta}$, $\tau$, and $\rho$, the definitions in (13) and (14), and the jump equations/inclusions of $x^{\circ}$, $\tau$, and $\rho$, will remain unchanged. Yet, under the jump equations of (53) and (54), the jump equations of $\tilde{\eta}_{p}$ and $\tilde{\zeta}_{pr}$ do change.
@@
Let $\mathcal{D}\subseteq\mathbb{R}^{mN}\times\mathbb{R}^{mNM^{\ast}}$, and consider the hybrid system $\mathcal{K}$ with flow set $\widetilde{C}\coloneqq C\times\mathcal{D}$, jump set $\widetilde{D}\coloneqq D\times\mathcal{D}$, single-valued flow map $\tilde{f}\colon\mathcal{X}\times\mathcal{D}\to\mathcal{X}$ by $\tilde{f}(\xi,\delta)\coloneqq f(\xi)$, and set-valued jump map $\widetilde{G}\colon\mathcal{X}\times\mathcal{D}\rightrightarrows\mathcal{X}$. Furthermore, let $\delta^{r}\coloneqq(\delta_{pr})_{p\in\mathcal{V}}$, $\delta\coloneqq((\delta_{p})_{p\in\mathcal{V}},(\delta^{r})_{r\in[M^{\ast}]})\in\mathcal{D}$,
$\widetilde{G}(\xi,\delta)$ $\coloneqq\{\widetilde{G}_{\tau,p}(\xi,\delta)\colon(\xi,\delta)\in D_{\tau,p}\times\mathcal{D}\text{ for }p\in\mathcal{V}\}\cup\{\widetilde{G}_{\rho,r}(\xi,\delta)\colon(\xi,\delta)\in D_{\rho,r}\times\mathcal{D}\text{ for }r\in[M^{\ast}]\},$ (55)
$\widetilde{G}_{\tau,p}(\xi,\delta)$ $\coloneqq\left[\begin{array}[]{c}x^{\circ}\\ \left[\tilde{\eta}_{1}^{\top},...,\tilde{\eta}_{p-1}^{\top},\delta_{p}^{\top},\tilde{\eta}_{p+1}^{\top},...,\tilde{\eta}_{N}^{\top}\right]^{\top}\\ \tilde{\zeta}\\ \left[\tau_{1},...,\tau_{p-1},[T_{1}^{p},T_{2}^{p}],\tau_{p+1},...,\tau_{N}\right]^{\top}\\ \rho\end{array}\right],\quad\widetilde{G}_{\rho,r}(\xi,\delta)\coloneqq\left[\begin{array}[]{c}x^{\circ}\\ \tilde{\eta}\\ \left[\tilde{\zeta}_{1}^{\top},...,\tilde{\zeta}_{r-1}^{\top},\big(\tilde{\zeta}_{pr}^{+}\big)_{p\in\mathcal{V}}^{\top},\tilde{\zeta}_{r+1}^{\top},...,\tilde{\zeta}_{M^{\ast}}^{\top}\right]^{\top}\\ \tau\\ \left[\rho_{1},...,\rho_{r-1},[T_{3}^{r},T_{4}^{r}],\rho_{r+1},...,\rho_{M^{\ast}}\right]^{\top}\end{array}\right],$
$\tilde{\zeta}_{pr}^{+}$ $=\left\{\begin{aligned} &\delta_{pr},&p\in\mathcal{V}^{r}\phantom{.}\\ &0_{m},&p\notin\mathcal{V}^{r}.\end{aligned}\right$.
Theorem 2.
@@
where $N^{\ast}\coloneqq N+M^{\ast}\in\mathbb{Z}_{>0}$, $p_{\max}\coloneqq\lambda_{\max}(\text{diag}(P_{2}(T_{2}),P_{3}(T_{4})))\in\mathbb{R}_{>0}$, and [^11] The matrices $P_{2}(\tau)$ and $P_{3}(\rho)$ are defined in (31), while the definitions of the constant vectors $T_{2}$ and $T_{4}$ are located above (31).
$\kappa_{2}\coloneqq\sqrt{\frac{p_{\max}N^{\ast}(2-\exp(-\mu T_{\min}/\alpha_{2}))}{\alpha_{1}(1-\exp(-\mu T_{\min}/\alpha_{2}))}}\in\mathbb{R}_{>0}.$
@@
$=\delta_{q}^{\top}P_{2,q}\exp(\sigma\tau_{q}^{+})\delta_{q}+(\delta^{s})^{\top}P_{3,s}\exp(\sigma\rho_{s}^{+})\delta^{s}-\tilde{\eta}_{q}^{\top}P_{2,q}\tilde{\eta}_{q}-\tilde{\zeta}_{s}^{\top}P_{3,s}\tilde{\zeta}_{s}$
@@
$\Delta V(\xi)\leq((\delta_{p})_{p\in\mathcal{V}})^{\top}P_{2}(T_{2})(\delta_{p})_{p\in\mathcal{V}}+((\delta^{r})_{r\in[M^{\ast}]})^{\top}P_{3}(T_{4})(\delta^{r})_{r\in[M^{\ast}]}\leq p_{\max}|\delta|_{\infty}^{2}.$ (58)
@@
$\text{dom}(\phi,\delta)\bigcap([0,t_{j}]\times\{0,1,...,j-1\})=\bigcup_{k=1}^{j}([t_{k-1},t_{k}]\times\{k-1\}).$
@@
$V(\phi(t,j))\leq\exp\left(-\frac{\mu}{\alpha_{2}}t\right)V(\phi(0,0))+p_{\max}|\delta|_{\infty}^{2}\sum_{s=1}^{j}\exp\left(-\frac{\mu}{\alpha_{2}}(t_{j}-t_{s})\right),$ (62)
@@
Let $j^{\ast}\coloneqq\lfloor j/N^{\ast}\rfloor\in\mathbb{Z}_{\geq 0}$. Note $j^{\ast}N^{\ast}\leq j$ and $j-j^{\ast}N^{\ast}\leq N^{\ast}$. The summation in (62) can be rewritten as
$\sum_{s=1}^{j}\exp\left(-\frac{\mu}{\alpha_{2}}(t_{j}-t_{s})\right)=\sum_{s=0}^{j^{\ast}-1}\sum_{k=1}^{N^{\ast}}\exp\left(-\frac{\mu}{\alpha_{2}}(t_{j}-t_{sN^{\ast}+k})\right)+\sum_{s=j^{\ast}N^{\ast}+1}^{j}\exp\left(-\frac{\mu}{\alpha_{2}}(t_{j}-t_{s})\right).$ (63)
@@
where $t_{j}^{s+1}\coloneqq t_{j}-t_{(s+1)N^{\ast}}$. Furthermore, since $\exp(-\tfrac{\mu}{\alpha_{2}}(t_{j}-t_{s}))\leq 1$ provided $t_{j}-t_{s}\geq 0$ and $t_{j^{\ast}N^{\ast}+1}\leq t_{j^{\ast}N^{\ast}+2}\leq...\leq t_{j}$, one has that
@@
$\sum_{s=0}^{\infty}N^{\ast}\exp\left(-s\frac{\mu}{\alpha_{2}}T_{\min}\right)=N^{\ast}\sum_{s=0}^{\infty}\exp\left(-\frac{\mu}{\alpha_{2}}T_{\min}\right)^{s}=\frac{N^{\ast}}{1-\exp(-\mu T_{\min}/\alpha_{2})}.$ (68)
@@
$V(\phi(t,j))\leq\exp\left(-\frac{\mu}{\alpha_{2}}t\right)V(\phi(0,0))+p_{\max}|\delta|_{\infty}^{2}N^{\ast}\frac{2-\exp(-\mu T_{\min}/\alpha_{2})}{1-\exp(-\mu T_{\min}/\alpha_{2})}.$ (70)
@@
7 Simulation Example & Discussion
@@
For each $p\in\mathcal{V}$, suppose the center of mass (CM) of agent $p$ is denoted by $c_{p}\coloneqq(c_{1,p},c_{2,p})\in\mathbb{R}^{2}$, expressed in a common inertial reference frame, and fixed relative to the body of the craft. The heading and angular velocity are denoted by $(\theta_{p},\omega_{p})\in(-\pi,\pi]\times\mathbb{R}$, and the linear speed is denoted by $v_{p}\in\mathbb{R}$. The control input of agent $p$ is $\mu_{p}\coloneqq(v_{p},\omega_{p})\in\mathbb{R}^{2}$.
For each marine craft $p\in\mathcal{V}$, let $\ell\in\mathbb{R}_{>0}$ denote the length between the CM and bow. Using the CM position $c_{p}$, heading $\theta_{p}$, and length $\ell$, the position of the bow of agent $p$ is given by $b_{p}\coloneqq c_{p}+\ell(\cos(\theta_{p}),\sin(\theta_{p}))$. Figure 2 provides a clarifying illustration. The goal of this simulation example is to achieve rendezvous in the bow positions, that is, $b_{p}=b_{q}$ for all $p,q\in\mathcal{V}$. To enable the use of our consensus result, we introduce a transformation. Utilizing (71), the time derivative of $b_{p}$ is
$\dot{b}_{p}=\underbrace{\begin{bmatrix}\cos(\theta_{p})&-\sin(\theta_{p})\\ \sin(\theta_{p})&\phantom{-}\cos(\theta_{p})\end{bmatrix}}_{\coloneqq R(\theta_{p})}\underbrace{\begin{bmatrix}1&0\\ 0&\ell\end{bmatrix}}_{\coloneqq D}\mu_{p}.$ (72)
Since $R(\theta_{p})$ is a rotation matrix and $\ell>0$, $R(\theta_{p})^{-1}$ and $D^{-1}$ are always well-defined. Furthermore, for any velocity $\dot{b}_{p}$, the corresponding control input is given by $\mu_{p}=D^{-1}R(\theta_{p})^{-1}\dot{b}_{p}$. Given a continuously differentiable reference trajectory for the bow position $b_{p}$, one can compute the corresponding control input for (71), pointwise in time, by utilizing the inverse map $\mathsf{T}^{-1}(\theta_{p},s)\coloneqq D^{-1}R(\theta_{p})^{-1}s$. Next, we provide a method for producing a reference trajectory for each $b_{p}$.
@@
where $z_{p}\in\mathbb{R}^{2}$ denotes a virtual position, $u_{p}\in\mathbb{R}^{2}$ is a virtual control input, $\mathtt{M}\in\mathbb{R}^{2\times 2}$ is a positive definite mass matrix, and $\mathtt{C}\in\mathbb{R}^{2\times 2}$ is a non-zero viscous damping matrix. The second-order ODE in (73) can be expressed as an LTI system by using $x_{1,p}\coloneqq z_{p}\in\mathbb{R}^{2}$, $x_{2,p}\coloneqq\dot{z}_{p}\in\mathbb{R}^{2}$, $x_{p}\coloneqq(x_{1,p},x_{2,p})\in\mathbb{R}^{4}$ using the notation from Section 2.1, and
$A\coloneqq\begin{bmatrix}0_{2\times 2}&I_{2}\\ 0_{2\times 2}&-\mathtt{M}^{-1}\mathtt{C}\end{bmatrix}\in\mathbb{R}^{4\times 4},\ B\coloneqq\begin{bmatrix}0_{2\times 2}\\ \mathtt{M}^{-1}\end{bmatrix}\in\mathbb{R}^{4\times 2},\ H\coloneqq\begin{bmatrix}I_{2}&0_{2\times 2}\end{bmatrix}\in\mathbb{R}^{2\times 4}.$ (74)
@@
By driving the auxiliary variables in $\{x_{p}\}_{p\in\mathcal{V}}$ into consensus, which will be achieved using the result from Section 4 and applying the transformation $\mathsf{T}^{-1}$ pointwise in continuous-time for each agent $p\in\mathcal{V}$, the marine crafts can accomplish rendezvous with respect to the bow positions. Note, however, our result does not guarantee collision avoidance between the marine crafts. Therefore, the trajectories generated by this approach can be treated as references to be tracked by a controller capable of ensuring collision avoidance. Two simulation cases are considered, namely, a nominal simulation that is free of perturbations and a more realistic simulation that contains measurement noise. The simulation parameters are the following: $N=14$, $n=4$, $d=m=2$, $M=5$, $M^{\ast}=8$, $\ell=0.53$, $\sigma=500$, $\mathtt{M}=I_{2}$, $\mathtt{C}=2.66\cdot I_{2}$, $K_{u}=0.5\cdot I_{2}$, and, for all $p\in\mathcal{V}$ and $r\in[M^{\ast}]$, $T_{1}^{p}=T_{3}^{r}=0.001$, $T_{2}^{p}=T_{4}^{r}=0.01$,
@@
Moreover, $x_{1,p}(0,0)=b_{p}(0,0)$ for each $p\in\mathcal{V}$. Thus, there are $14$ agents in total that are divided into $5$ clusters and $8$ inter-clusters. The parameters $n$, $d$, and $m$ are determined by the LTI system corresponding to the auxiliary system in (73). The length $\ell$ between the bow and CM of each marine craft was chosen arbitrarily. Likewise, the parameters $\sigma$, $T_{1}^{p}$, $T_{3}^{r}$, $T_{2}^{p}$, and $T_{4}^{r}$ and matrices $\mathtt{M}$, $\mathtt{C}$, $K_{u}$, $K_{\eta,p}$, and $K_{\zeta,pr}$ were chose at will provided $0<T_{1}^{p}\leq T_{2}^{p}$ for each agent $p\in\mathcal{V}$, $0<T_{3}^{r}\leq T_{4}^{r}$ for each inter-cluster $r\in[M^{\ast}]$, and the matrix inequality in (32) could be satisfied at $(\tau,\rho)=(0_{N},0_{M^{\ast}})$ and $(\tau,\rho)=(T_{2},T_{4})$. Under these parameters, MATLAB and the CVX Toolbox [Grant.Boyd2008, Grant.Boyd2014] were employed to solve the matrix inequality $\mathsf{M}(\tau,\rho)\leq\mathbf{0}$ for points $(\tau,\rho)=(0_{N},0_{M^{\ast}})$ and $(\tau,\rho)=(T_{2},T_{4})$. Consequently, Corollary 1 implies the sufficient conditions of Theorem 1 are satisfied. Unfortunately, we are unable to provide the solution to the matrix inequality since $P(0_{N},0_{M^{\ast}})\in\mathbb{R}^{304\times 304}$ in our application. Note, CVX took about 100 seconds to compute $P(0_{N},0_{M^{\ast}})$.
The parameters and communication topology, i.e., the clustered network depicted in Figure 3, are identical for the nominal and perturbed simulation cases. Moreover, the initial condition of each agent was selected in a manner leading to the construction of a distance-dependent graph with nodes in $\mathbb{R}^{4}$ having edges no longer than $25$ units of length. Figures 46 correspond to the simulation results for the nominal case, while Figures 78 correspond to the simulation results for the perturbed case. To examine the robustness of our consensus controller to various perturbation magnitudes and communication graphs, a modest Monte Carlo simulation was performed with results provided in Figure 9. Note only the graph $\mathcal{G}$, perturbation magnitudes, and controller matrix $K_{u}$ were changed in the Monte Carlo simulation. The graphs were defined in a similar fashion to the nominal and perturbation simulation cases described above, leading to corresponding initial conditions for each agent. The controller matrix $K_{u}=1.5\cdot I_{2}$ was fixed for all $30$ runs of the Monte Carlo simulation. However, perturbation magnitudes were randomly selected from the interval $[0,0.075]$ using a uniform distribution. The partition of each graph’s node set $\mathcal{V}$ was also randomly selected.
@@
7.1 Consensus under Ideal Conditions
As validated by Figures 46, the proposed control strategy in Section 4 renders the set $\mathcal{A}$ GES under ideal conditions (i.e., under the absence of measurement noise). Consequently, the reference trajectories for the marine crafts can be synchronized exponentially fast while using intermittent and asynchronous output feedback (that is, position measurements). Also, the map $(\theta_{p},x_{2,p})\mapsto D^{-1}R(\theta_{p})^{-1}x_{2,p}$ enabled each agent $p\in\mathcal{V}$ to track their reference trajectory. Recall $x_{2,p}$ denotes the velocity of the reference trajectory for agent $p$. Therefore, we were able to achieve exponentially fast rendezvous for marine crafts modeled by (71).
Of course, driving $|\xi|_{\mathcal{A}}$ to 0 in practice would probably be undesirable as the user may not wish for potentially expensive marine crafts to crash into each other. Thus, achieving $|\xi|_{\mathcal{A}}\leq\omega$ for some appropriate $\omega>0$ may suffice instead. Given such a scenario, $|\xi|_{\mathcal{A}}\leq\omega$, (35), and (42) facilitate the calculation of a conservative, albeit finite, continuous time for achieving $\omega$-approximate rendezvous (see [Zegers.Guralnik.ea2021, Defintion 2]). In particular,
$t^{\ast}\coloneqq\frac{2\alpha_{2}}{\mu}\ln\left(\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}\frac{|\phi(0,0)|_{\mathcal{A}}}{\omega}\right)\leq t\text{ and }0<\omega<\sqrt{\frac{\alpha_{2}}{\alpha_{1}}}|\phi(0,0)|_{\mathcal{A}}\implies|\phi(t,j)|_{\mathcal{A}}\leq\omega$